OpenStudy (anonymous):

Which of the following equations best represents the graph of a function f(m)? (1 point) f(y) = m f(x) = m y = f(m) x = f(m)

OpenStudy (anonymous):

Pleas help will fan and medal the person who explains it and gives me the answer!

OpenStudy (anonymous):

@smilobunny can help you out.

OpenStudy (anonymous):

y = f(m)

OpenStudy (anonymous):

y = a function of m

OpenStudy (anonymous):

Ok thanks can you help me with some more?

OpenStudy (anonymous):

sure one more

OpenStudy (anonymous):

If f(x) = 7x - 1, what does f(12) represent? The value if x when (7x - 1) = 12 The value of (7x - 1) when x = 12 The value of 12 f(x) when x = 12 The value of f(x) when (7x - 1) = 12

OpenStudy (anonymous):

plug in 12 for x

OpenStudy (anonymous):

The value of (7x - 1) when x = 12

OpenStudy (anonymous):

So B?

OpenStudy (anonymous):

f(x) F(12) x=12

OpenStudy (anonymous):

@Deeezzzz please do not just solely give answers please. This is against the guidelines.

OpenStudy (anonymous):

But deeezzzz is correct

OpenStudy (anonymous):

If f(x) = 2x2 - 10, find f(5). 90 10 7.5 40 Last one haha

OpenStudy (anonymous):

Plug in 5 for x

OpenStudy (anonymous):

Mkay...

OpenStudy (anonymous):

f(5) is essentially f(5)=(2x5)^2 -10

OpenStudy (anonymous):

Mkay.... so it would be C?

OpenStudy (anonymous):

uh....nooo... Is that a 2x^2 btw? just checking

OpenStudy (anonymous):

2x with a 2 has the exponent

OpenStudy (anonymous):

so \[f(x)=2x^2-10\]

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

then \[f(5)=(2\times5)^2-10\]

OpenStudy (anonymous):

Ohh wait a moment would it be A?

OpenStudy (anonymous):

\[f(5)=(10)^2-10\]

OpenStudy (anonymous):

YES Correct!!!

OpenStudy (anonymous):

Ok thanks so much!!!

OpenStudy (anonymous):

np, if you need any more help just tag me

OpenStudy (anonymous):

If f(x) is the height, in cm, of a sunflower plant which is x days old, which of the following statements best describes the meaning of f(60) = 210? The height of the sunflower plant is 60 cm when it is 210 days old. The height of the sunflower plant is 210 cm when it is 60 days old. The height of the sunflower plant is 210 cm when it is 3.5 days old. The height of the sunflower plant is 60 cm when it is 3.5 days old.

OpenStudy (anonymous):

Abby, do another post, this is getting very lengthy

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol good idea